OpenStudy (anonymous):

solving exponential equations solve for x: 8 ^x+2= 16^x

OpenStudy (anonymous):

how too find the answer

OpenStudy (anonymous):

taking logs (x+2) ln 8 = x ln 16 (x+2) / x =ln 16 / ln8 = 4/3 3x + 6 = 4x x = 6

OpenStudy (anonymous):

get the 8 and 16 to become the same base... so 8 can be 2^3 ... and 16 can be 2^4 so right now... (2^3)^(x+2) = (2^4)^x since you have the same base, cancel them out you now use the exponents as your equation... 3(x+2) = 4(x) 3x+2=4x solve and you get x=6

OpenStudy (anonymous):

yep - a good answer wendy